OpenStudy (anonymous):

whats the highest common factor of 630 and 1560?

OpenStudy (anonymous):

"To find the GCF of two numbers: List the prime factors of each number. Multiply those factors both numbers have in common. If there are no common prime factors, the GCF is 1." - Google.

OpenStudy (gorv):

first factorise them sepratelly

OpenStudy (gorv):

thn take the factor which are common in both

OpenStudy (gorv):

welcome to OS :)

OpenStudy (gorv):

u know how to factor ??

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

can you finish the prime factorisation of 1560 @Sarmster ?

OpenStudy (anonymous):

I worked out most of it just not sure what to do now 3*3*7*5*2 = 630 3*5*2*2*3*3*3= 1560 3*3*2*7 = 126 (HCF) Is this correct

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

the first one is right , but 3*5*2*2*3*3*3 = 1620, what steps did you have for factoring 1560?

OpenStudy (unklerhaukus):

you should get a factor of 13